a) Find change in wavelength of 80-ppm x-rays that are scattered 120 degree by a target.

b)Find the angle between the directions of recoil electron and incident photon.

c) Find the energy of recoil electron.


1
Expert's answer
2022-03-02T17:00:00-0500

"\Delta \lambda=\lambda(1-\cos\varphi)=3.6~pm,"

"\theta=\arctan(1-\frac{\Delta \lambda}{\lambda})=87.4°,"

"E=hc(\frac 1{\lambda}-\frac 1{\lambda+\Delta\lambda})=9.3~\cdot 10^{-17}~J ."